OpenStudy (anonymous):

What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (3, 6) and (1, −2)?

OpenStudy (anonymous):

First find the slope by using the slope formula \[ m = \frac{Y_2 - Y_1}{X_2 - X_1} \] Just plugin your two point values (3, 6) and (1, −2) \[ m = \frac{(-2) - 6}{1 - 3} = \frac{-8}{-2} = 4\] Now that we know our slope, which is 4, we have our answer because they want the slope of a line that is parallel. Parallel lines have the same slope. So the slope of a line that passes through the point (−1, 1) is 4